OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 - 10x2 + 24x ? -6, 0, 4 6, 0, -4 -6, 0, -4 6, 0, 4

OpenStudy (anonymous):

Factor, then use the zero product property. Remember that the first factoring step is to remove the common factors (in this case, x).

OpenStudy (anonymous):

\[f(x)=x^3−10x^2+24x=0=x(x−6)(x−4) \implies x∈\left\{ 0,4,6 \right\}\]